# Index Terms:

### v1

@@ VII-B System Performance Comparison
  • DBAG [20]: A state-of-art (SOTA) DRL method that formulates problem (13) as an MDP and applies soft actor-critic for degradation action optimization.
@@
Following discrete degradation action selection, all benchmark methods derive optimal bandwidth allocations by solving subproblem (27). While commercial CVX solvers such as MOSEK [30] guarantee optimal solutions for (27), they incur substantial computational overhead when deployed on an ES equipped with low-frequency CPU. This limitation is especially pronounced for the IDEAL method, which requires solving (27) for $A^{N}$ times per slot due to exhaustive search. To mitigate computational complexity, we implement a DNN-based bandwidth allocation solver following [20], which efficiently approximates solutions to (27). Fig. 6 compares the performance of MOSEK and our DNN solver when implementing the IDEAL method over 1,000 time slots. The results demonstrate that the DNN solver achieves almost identical per-slot utility to the optimal solution. Based on this validation, we employ the DNN solver for all subsequent method evaluations. We conduct Monte Carlo simulations across all methods using 10 random seeds. Each simulation spans $T=3,000$ time slots, with final results averaged over all seeds.
@@ VIII Conclusion
Appendix A Definitions of the Constituent Kernels
To effectively model the composite structure of the input vector $\bm{z}_{t}=\left(\bm{h}_{t},\bm{a}_{t},t\right)$, we consider three specialized kernels to capture the distinct nature of each component. In the following, the subscript $i$ and $i^{\prime}$ denote time indices.
1) Channel State Kernel: To capture the similarity in continuous channel states $\bm{h}_{t}$, we utilize an automatic relevance determination (ARD) radial basis function (RBF) kernel:
2) Degradation Action Kernel: To handle discrete degradation actions $\bm{a}_{t}$, we implement a categorical kernel [31]:
3) Temporal Dynamics Kernel: To model time-varying correlations arising from evolving channel conditions and content characteristics, we employ a temporal kernel [31]:
Appendix B Proof of Proposition 1
[…]
By setting (46) to 0, we obtain
$\tau_{t,n}^{\rm o\ast}=\sqrt{\frac{\phi_{n}^{\ast}d_{t,n}}{w_{n}}}.$ (48)
By setting (47) to 0, we have
By solving (49), we obtain that
where $\mathcal{W}(\cdot)$ is the Lambert-$\mathcal{W}$ function.

### v2

@@ VII-C System Performance Comparison
  • DBAG [23]: A SOTA DRL method that formulates problem (13) as an MDP and applies soft actor-critic for degradation action optimization.
@@
For a fair comparison, all training-based methods (BO, DBAG, and our proposed LAB) are evaluated under an online training paradigm without using any pre-trained models. Besides, we strictly adhere to the original neural network architecture of DBAG [23] for its actor and critic networks. Following discrete degradation action selection, all benchmark methods derive optimal bandwidth allocations by solving subproblem (30). While commercial CVX solvers such as MOSEK [34] guarantee optimal solutions for (30), they incur substantial computational overhead when deployed on an ES equipped with low-frequency CPU. This limitation is especially pronounced for the IDEAL method, which requires solving (30) for $A^{N}$ times per slot due to exhaustive search. To mitigate computational complexity, we implement a DNN-based bandwidth allocation solver following [23], which efficiently approximates solutions to (30). Fig. 7 compares the performance of MOSEK and our DNN solver when implementing the IDEAL method over 1,000 time slots. The results demonstrate that the DNN solver achieves almost identical per-slot utility to the optimal solution. Based on this validation, we employ the DNN solver for all subsequent method evaluations. We conduct Monte Carlo simulations across all methods using 10 random seeds. Each simulation spans $T=3,000$ time slots, with final results averaged over all seeds.
@@ VIII Conclusion and Discussions
Figure 11: Performance comparison of LAB and baseline methods on VisDrone2019 dataset.
We conclude this paper with an extended evaluation of our LAB framework on the real-world VisDrone2019 dataset [35]. In this simulation, we set $\omega=4$ and $\zeta=\sqrt{2}$. Fig. 11 plots the moving average of the utility $U_{t}$ over the last 500 slots. The results show that LAB consistently outperforms all baseline methods in this new scenario, improving the utility in average by 11.33%, 27.98%, 90.91%, and 167.30%, respectively. This outcome supports the potential generalizability of LAB and encourage its further exploration in diverse edge‑AI settings. Notably, the high scene complexity and object density of VisDrone2019 results in a 23.95% performance gap between LAB and the IDEAL upper bound. This highlights a clear direction for future work, where further gains could be achieved through specialized optimizations in kernel design, DNN architecture, and hyper‑parameters.